Sheldon hires Raj for $5 and sells daisies to Leonard for $15. Leonard rents space from Amy for $5 and sells bouquet to Penny for $30. Penny gives ECON 102 lectures to Sheldon for $10. What is GDP?

The answer is 40.


I don’t understand why GDP (from production perspective): 10+30=40

Why don’t we count $5 that’s paid to Raj or $5 rental fee?

Also, if I calculated GDP from income perspective, our teacher told us that Leonard’s income is 30-5-15=10 and Sheldon’s income is 15-5=10

Why do we need to subtract 15 from 30 for Leonard?
If we need to use revenue minus expenditure, then for the case of Sheldon, why don’t we also subtract the lecture fee from his total earning of $15?

I’m really confused, thank you for anyone who can help me with it!
Answer #1

1. From production perspective, we only add the value of final product in the calculation of GDP. Here, the final product is the bouquet and the lectures by Leonard. Thus, GDP = 10+30=40.

2. Leonard purchased two inputs- i) daisies from Sheldon for $15 and ii) rent space from Amy. Thus, Leonard's net income after accounting for these inputs- 30-15-5 = 10.

Sheldon's income is 15-5= 10.

Amy's income is 5 and Raj's income is also 5.

Hence, from income perspective GDP = 10+10+5+5 = 40.0

3. It's actually revenue minus expenditure from the production process. The lectures taken by Sheldon is his consumption expenditure and hence will not be considered if we calculate GDP by income or production method. It will rather be considered as a consumption expenditure if we calculate GDP by expenditure method.

Hope this helps.

Thanks!